OpenStudy (elenathehomeschooler):

8. A geometric sequence has a4=4 and a5=7 . What is a1 ?

OpenStudy (elenathehomeschooler):

Can you alos help me with this? @jim_thompson5910 I understand how to find partial sums but i dont know how to find a1

jimthompson5910 (jim_thompson5910):

what is the common ratio here? are you able to find it?

OpenStudy (elenathehomeschooler):

its 1.75

jimthompson5910 (jim_thompson5910):

correct, so r = 1.75

jimthompson5910 (jim_thompson5910):

the nth term of a geometric sequence is \[\Large a_{n} = a_1*(r)^{n-1}\]

jimthompson5910 (jim_thompson5910):

we're given \[\Large a_4 = 4\] so n = 4 and \(\Large a_n = 4\) \[\Large a_{n} = a_1*(r)^{n-1}\] \[\Large a_{4} = a_1*(1.75)^{4-1}\] \[\Large 4 = a_1*(1.75)^{4-1}\] are you able to isolate \(\Large a_1\) ?

OpenStudy (elenathehomeschooler):

yes

OpenStudy (elenathehomeschooler):

would a1= 1.3593

jimthompson5910 (jim_thompson5910):

I think you divided in the wrong order

jimthompson5910 (jim_thompson5910):

I'm getting 0.7463556851312

jimthompson5910 (jim_thompson5910):

if you kept everyhing as a fraction, then \[\Large n = 4\] \[\Large a_n = 4\] \[\Large r = \frac{7}{4}\] would lead to \[\Large a_1 = \frac{256}{343}\] Notice how \[\Large \frac{256}{343} \approx 0.7463556851312\]